OpenStudy (rabya!):

(5e^x)^3 , how can i solve this? answer is 375e3x

OpenStudy (rabya!):

Its deerivative >

OpenStudy (rabya!):

derivative*

OpenStudy (anonymous):

\[y= [f(x)]^n\] \[y'=n \times [f(x)]^{n-1} \times f'(x)\]

OpenStudy (anonymous):

\[y=(5e^x)^3\] \[y'=3(5e^x)^2 \times 5e^x\] \[y'=3 \times 25 \times e^{2x} \times 5 \times e^x\] \[y'=375e^{3x}\]

OpenStudy (rabya!):

Thank You Ankh ! =)